A 74.8 g sample of a metal at 143.2 °C is added to an insulated vessel containing

208g of water at 24.8 °C. The final temperature of the mixture is 31.1 °C. What is

the specific heat of the metal in J/g-°C?


1
Expert's answer
2022-02-18T05:29:01-0500

mass of metal= 74.8g

temperature of metal= 143.2 degrees Celsius

mass of water= 208g

temperature of water= 24.8 degrees Celsius

Final temperature of mixture=31.1 degrees Celsius

Specific heat of metal × mass of metal × (final temperature of mixture - temperature of metal) = Specific heat of water × mass of water × (final temperature of mixture - temperature of water)

Specific heat of metal × 74.8 ×(-112.1) = 4.184 × 208 × (6.3)

Specific heat of metal= (5482.7136 ÷ 8385.08) j/g degrees celsius

Specific heat of metal :

= 0.654 J/g degrees celsius